The stock of Simon will go ex-dividend tomorrow. The dividend will be $1 per share, and there are 15,000 shares of stock outstanding. The market-value balance sheet for Simon is shown below.

                   

Assets Liabilities and Equity
Cash $250,000      Equity $1,000,000     
Fixed assets 750,000     

  

So far, price of the share today is $66.67 per share and it will sell at $65.67 per share for tomorrow.
Now suppose that Simon announces its intention to repurchase $15,000 worth of stock instead of paying out the dividend. (Round your answers to the nearest cent.)

               

a. What effect will the repurchase have on an investor who currently holds 200 shares and sells 3 of those shares back to the company in the repurchase?

    

Total value of the position $

       

b. What effect will the initial cash dividend payment have on the same investor?

    

Total value of the position $
Answer #1
a) Market value of assets after repurchase = 1000000-15000 = $    9,85,000
Number of shares outstanding after repurchase = 15000-15000/66.67 = 14775
Share price after repurchase = 985000/14775 = $          66.67
Effect on the investor:
Total wealth before repurchase = 200*66.67 = $        13,334
Total wealth after repurchase:
Value of shares on hand = 197*66.67 = $        13,134
Cash on sale back = 3*66.67 = $              200
Total wealth after repurchase $        13,334
There will be no effect on the shareholders wealth.
b) Total wealth before dividend payment = 200*66.67 = $        13,334
Total wealth after dividend payment:
Market value of shares = 200*965.67 = $        13,134
Cash received for dividends = 200*1 = $              200
Total wealth after dividend payment $        13,334
There will be no effect on the shareholders wealth.